# Bound-electron self-energy calculations in Feynman and Coulomb gauges: detailed analysis

### v1

@@ Abstract
The energy correction associated with the self-energy diagram is the leading (in magnitude) contribution to the Lamb shift in hydrogen-like ions. All conventional approaches to this correction rely on partial-wave expansions, which are a stumbling block limiting accuracy. To elucidate an issue, we perform a comparative analysis of partial-wave-expansion convergence in two gauges: Feynman and Coulomb. Some tricks for improving convergence are also discussed.
@@ I Introduction
Accurate theoretical predictions for simple atomic systems must include QED effects mohr_main; indelicato_qed_2019; SHABAEV202494; volotka_progress_2013. One of the leading QED corrections to the bound-electron energy levels corresponds to the one-electron self-energy (SE) diagram, shown in Fig. 1. A high-precision evaluation of the corresponding contribution holds significance, however, not only for studies of the Lamb shift and transition energies in hydrogen-like ions and other few-electron systems Stohlker; PhysRevLett.94.223001; PhysRevLett.66.1434; PhysRevLett.91.073202; PhysRevLett.95.233003; Trassinelli_2009; Trassinelli_2011; Loetzsch2024 (for related theory, see, e.g., Refs. PhysRevLett.97.253004; PhysRevA.77.032501; PhysRevA.83.012504; se_yerokh; PhysRevA.71.062104; PhysRevA.100.062506), but also for consideration of a wide range of radiative corrections. Accurate calculations of this contribution is a cornerstone in the ab initio treatment of the two-electron SE diagrams PhysRevLett.76.204; PhysRevA.60.3522; PhysRevA.63.052507, QED corrections to the $g$-factor yerokhin_self-energy_2002; PhysRevA.62.032510, quadratic Zeeman splitting agababaev2025qedeffectsquadraticzeeman, and hyperfine structure PhysRevA.64.012506; PhysRevA.63.032506; PhysRevLett.100.163001; PhysRevA.81.012502. In addition, the SE correction is essential when studying the E1 E1_Kozlov and M1 M1_Volotka transition amplitudes. In these applications it is often necessary to calculate off-diagonal matrix elements of the SE operator. However, one-electron (diagonal) SE is a simple example that one can and should use to experiment in order to understand how to calculate it with the highest precision.
There are two fundamentally different approaches to the SE calculations. The first one is based on the nonrelativistic formulation of QED and implies an expansion in $\alpha Z$ caswell ($\alpha$ is the fine-structure constant and $Z$ is the nuclear charge). It is suitable for light nuclei, in which this parameter is much less than unity. In contrast, the second approach is non-perturbative in $\alpha Z$, i.e., all calculations are carried out to all orders in this parameter, see, e.g., Refs. SHABAEV2002119; mohr_main. These approaches can be applied as complementary RevModPhys.97.025002. In the present paper, we employ the second approach, also known as the Furry picture of QED PhysRev.81.115, to calculate the SE correction.
@@
In this regard, an issue of representing the bound-electron propagator in calculations naturally arises. Unlike the free-electron propagator, the propagator in a binding potential does not have a closed-form expression. However, when the spherical symmetry is present, as is the case in the atomic calculations, such the propagator can be constructed, e.g., using a partial-wave (PW) expansion. With this in mind, we use two state-of-the-art approaches: the Green’s function (GF) method MOHR197426; MOHR197452; PhysRevA.38.5066, and the finite-basis-set (FBS) method PhysRevA.37.307; DKB.
@@
Within the first approach, the bound-electron propagator is represented in terms of the Dirac-equation solutions bounded at infinity and at origin. For the point-nucleus Coulomb potential, the corresponding solutions are expressed analytically via the Whittaker functions PhysRev.101.843. For more realistic models of nuclear-charge distribution, e.g., the Fermi model, or other spherically symmetric potentials, e.g., some local screening potentials, these solutions can be found numerically. This approach enables one to perform calculations with a high numerical accuracy when large numerical cancellations occur or when the PW expansion does not converge rapidly. Note that the Green’s function, considered as a function of any of its radial arguments, is discontinuous, when the arguments coincide. Therefore, obtaining accurate results for the matrix elements of the Green’s function requires special care.
@@
Within the second approach, the electron Green’s function is built using the FBS method. Basis sets for the Dirac equation could be constructed, e.g., from the B splines J_Sapirstein_1996; johnson or Gaussians boys_gauss; se_gauss. The corresponding numerical procedures can easily incorporate any spherically symmetric potential. However, for systems that do not have spherical symmetry, the FBSs can also be readily prepared PhysRevA.106.012813; DIRAC25. The advantage of this method is that it allows one to easily separate out and exculde, if necessary, the contribution of a specific bound state to the Green’s function. Note also that this method provides an approximation to the Green’s function, which is a continuous function of the radial arguments. Nevertheless, the basis-set method has some important drawbacks when compared to the GF approach. Because of an additional parameter, the number of basis functions $N$, the corresponding results should be extrapolated to $N\rightarrow\infty$. This sets a limit on the computational accuracy. Furthermore, the number of partial waves, usually considered in practical calculations by means of the FBS method is limited and inferior to that in the Green’s function method sym12050800. In this paper we use the FBS method with the basis obtained from B splines within the double-kinetic-balance (DKB) approach DKB. The latter method excludes the so-called spurious states and establishes the correct asymptotics of wave functions in the non-relativistic limit. The effectiveness of this basis as applied to the calculations of the SE correction has been demonstrated, e.g., in Ref. DKB.
@@
The SE diagram contains ultraviolet (UV) divergences. The conventional approach to renormalizing these divergences, which is used in the present work, involves the potential expansion of the bound-electron propagator snyderman. Within this expansion, three contributions naturally arise: zero-, one-, and many-potential terms. Only the zero- and one-potential terms, which correspond to the two first terms of the potential expansion, are UV divergent. They are treated separately in momentum space. The many-potential term, which is given by a remainder of the potential expansion, is considered in coordinate space using the PW expansion. The truncation of the latter and subsequent estimation of the residual is the main source of numerical uncertainty of the result.
@@
As is well known, the total SE contribution is gauge invariant, unlike the individual terms of its potential expansion. Therefore, an intriguing idea is to compare the magnitude of these terms in different gauges. Traditionally, the SE correction is calculated in the Feynman gauge, see, e.g., Ref. base_1999. In the Coulomb gauge, similar calculations were carried out in hen; yer_2025. These works suggest that the many-potential contribution is significantly smaller in the Coulomb gauge than in the Feynman one. However, a smaller absolute value does not imply a better convergence of the PW expansions. Therefore, an important task is to conduct a thorough comparative analysis of many-potential contributions in both gauges. Note that other covariant gauges, aside from the Feynman one, can also be applied to the SE calculations. For example, the Fried-Yennie PhysRevD.47.3647 gauge was also considered in Ref. yer_2025. In the present work, however, we restrict our consideration only to the Feynman and Coulomb gauges.
@@
Another important task in the SE calculations is to somehow accelerate the convergence of employed PW expansions. It is reasonable to assume that the poor convergence of the many-potential contribution is due to the leading term of its potential expansion. Therefore, the first natural idea is to additionally subtract the two-potential term from it, which should improve the PW-expansion convergence of the remainder. The subtracted term has to be evaluated separately with high precision and added to the final result. In the following, we will refer to this approach as the two-potential scheme. It should be noted, however, that the direct calculation of the two-potential term in momentum space in a closed form, i.e., not resorting to the PW expansion, turns out to be a challenging problem PhysRevLett.100.163001; Yer2010.
@@
Another acceleration-convergence approach, which we will further refer to as the Sapirstein-Cheng (SC) scheme, was proposed recently in Ref. sap_orig. Within the framework of this promising scheme, an approximation to the two-potential term, the quasi-two-potential contribution, can be subtracted from the many-potential term instead. In contrast to the exact two-potential term, its quasi-two-potential counterpart can be easily calculated in momentum space. For the point-nucleus case, calculations of the SE correction using this acceleration scheme have been performed in Ref. yer_2025 for different gauges. However, it is also noteworthy to examine the case of an extended nuclues.
@@
We also note another acceleration-convergence approach developed earlier for the SE correction to the Lamb shift in Ref. PhysRevA.72.042502, which is referred to as the Yerokhin, Pachucki, and Shabaev (YPS) scheme. It also improves the convergence by subtracting an approximation for the many-potential contribution, that can be calculated in a closed form. The YPS scheme was initially created for the calculations in the Feynman gauge, and was later generalized to the Coulomb-gauge case in Ref. yer_2025. Judging by the results of the latter paper, it is an extremely powerful scheme, however, its drawback is that it is difficult to generalize it to more complex diagrams, unlike other acceleration-convergence approaches. For instance, the two-potential scheme has recently been applied in the QED calculations of the quadratic Zeeman effect agababaev2025qedeffectsquadraticzeeman, and some modifications of the SC scheme have been applied to the calculations of the two-electron SE sap_alexey and two-loop SE PhysRevLett.133.251803; PhysRevA.111.042820 contributions. Due to this fact, we do not consider the YPS scheme in the current study.
@@
In this paper, we thoroughly examine the application of several different schemes to the SE calculations in both the Coulomb and Feynman gauges. We conduct an independent of Ref. yer_2025 analysis, consider in details the issue of the PW-expansion convergence, perform the SE-correction calculation for nonpoint nuclei, and collect convenient formulas that allow one to make the corresponding calculations.
@@
Throughout this article the relativistic units ($\hbar=c=1$) are used. We use roman style ($\rm{p}$) for four-vectors, bold style ($\bf{p}$) for space vectors, and italic style ($p$) for scalars.
@@ II.1 Basic formalism
Within the framework of the Furry picture, the description of an atom or ion starts with the Dirac equation:
$\left[-i(\boldsymbol{\alpha}\cdot\boldsymbol{\nabla})+\beta m+V(\mathbf{r})\right]|a\rangle=\varepsilon_{a}|a\rangle,$ (1)
where $|a\rangle$ and $\varepsilon_{a}$ are the Dirac wave function and energy, $\boldsymbol{\alpha}$ and $\beta$ are the Dirac matrices, $V(\mathbf{r})$ is a binding potential, which is a function of the position vector $\mathbf{r}$. In the following, we consider only hydrogen-like ions for which the potential is spherically symmetric, $V(\mathbf{r})=V(r)$, $r=|\mathbf{r}|$. Then, the solution of Eq. (1) can be represented in the form:
@@
Figure 1: Self-energy diagram with the related mass counterterm. The double line indicates the electron propagator in the external field of the nucleus, the wavy line denotes the photon propagator, and the cross stands for the counterterm.
The energy shift $\Delta\varepsilon_{a}$ of the bound state $|a\rangle$ due to the first-order self-energy correction, which is graphically represented in Fig. 1, is given base_1999 by the real part of the expression:
@@
where $\mathbf{r}_{12}=\mathbf{r}_{1}-\mathbf{r}_{2}$, $D_{\mu\nu}(\omega,\mathbf{r}_{12})$ is the photon propagator, $\alpha_{\mu}=(1,{\boldsymbol{\alpha}})$, $G(\varepsilon,\mathbf{r}_{1},\mathbf{r}_{2})$ is the bound-electron Green’s function, and $\delta m$ is the mass counterterm.
Let us define the gauge-dependent operator $I(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ according to
@@
In the Feynman and Coulomb gauges, this operator has the forms $I_{F}$ and $I_{C}$, respectively:
$I_{F}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\alpha\left[1-({\boldsymbol{\alpha}}_{1}\cdot{\boldsymbol{\alpha}}_{2})\right]\frac{e^{i\hat{\omega}r_{12}}}{r_{12}},$ (5)
$I_{C}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\alpha\left[\frac{1}{r_{12}}-({\boldsymbol{\alpha}}_{1}\cdot{\boldsymbol{\alpha}}_{2})\frac{e^{i\hat{\omega}r_{12}}}{r_{12}}+\frac{(\boldsymbol{\alpha}_{1}\cdot\boldsymbol{\nabla}_{1})(\boldsymbol{\alpha}_{2}\cdot\boldsymbol{\nabla}_{2})}{\hat{\omega}^{2}}\frac{e^{i\hat{\omega}r_{12}}-1}{r_{12}}\right],$ (6)
where $r_{12}=|\mathbf{r}_{12}|$, $\hat{\omega}=\sqrt{\omega^{2}+i0}$, and the branch of the square root is fixed with the condition $\Im\sqrt{\omega^{2}+i0}>0$.
In its turn, the bound-electron propagator can be expressed using the spectral representation:
$G(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\sum_{n}\frac{\psi_{n}(\mathbf{r}_{1})\psi_{n}^{\dagger}(\mathbf{r}_{2})}{\omega-\varepsilon_{n}^{-}},$ (7)
where the summation runs over the complete Dirac spectrum and $\varepsilon_{n}^{-}=\varepsilon_{n}(1-i0)$. In shortened form, the formula (7) can be written as follows:
$G(\omega)=\sum_{n}\frac{|n\rangle\langle n|}{\omega-\varepsilon_{n}^{-}}.$ (8)
Then, the energy shift (3) can be expressed as:
@@
where the diagonal matrix element of the one-loop self-energy operator $\Sigma(E)$ is given by:
$\langle a|\gamma^{0}\Sigma(E)|a\rangle=\frac{i}{2\pi}\int_{-\infty}^{\infty}d\omega\sum_{n}\frac{\langle an|I(\omega)|na\rangle}{E-\omega-\varepsilon_{n}^{-}},$ (10)
and $\gamma^{0}\equiv\beta$. The expression (10) contains the UV divergences. To properly treat them, we expand the bound-electron propagator in powers of the binding potential blundell_snyderman. The potential expansion can conveniently be written as:
@@
where the states $|n\rangle$ correspond to the bound-electron spectrum, the states $|f\rangle$, $|f_{1}\rangle$, and $|f_{2}\rangle$ correspond to the free-electron spectra, and the last term is expressed via the bound-electron Green’s function $G(\omega)$ itself and encompasses an infinite number of terms corresponding to the two or more interactions with the nucleus. As before, all summations in Eq. (11), are carried out over the complete spectra for the corresponding Dirac hamiltonians.
Figure 2: Potential expansion of the self-energy diagram. The dashed line ended by a rhombus denotes the interaction with the nuclear field. The single line denotes the free-electron propagator. The mass counterterm is not shown.
Substituting the potential expansion (11) into Eq. (10) splits the SE diagram into three terms, as shown graphically in Fig. 2. These contributions are commonly referred to as the zero-, one-, and many-potential terms, respectively. The zero- and one-potential terms diverge, whereas the many-potential term is UV finite. The UV-divergent terms have to be renormalized along with the mass counterterm. For this aim, they are considered in momentum space, where the dimensional regularization is applied to isolate and eliminate the UV divergences. This renormalization procedure yields UV-finite contributions $\Delta\varepsilon_{a}^{\text{0p}}$ and $\Delta\varepsilon_{a}^{\text{1p}}$ to the Lamb shift. The many-potential term $\Delta\varepsilon_{a}^{\text{Mp}}$ is calculated in the coordinate space. Therefore, the SE correction to the energy levels is expressed as the gauge-invariant sum of three contributions:

### v2

@@ Abstract
The energy correction associated with the self-energy diagram is the leading (in magnitude) and fundamental (in significance) contribution to the Lamb shift in highly charged ions. Conventional approaches to this correction rely on partial-wave expansions, which is a stumbling block limiting accuracy. To elucidate the issue, we perform a comprehensive comparative analysis of partial-wave-expansion convergence between two gauges: Feynman and Coulomb. Some tricks for improving the convergence are discussed as well.
@@ I Introduction
Quantum electrodynamics (QED) effects are indispensable for the accurate theoretical description of various atomic systems. To account for relativistic effects, the corresponding calculations are to be performed non-perturbatively in the nuclear-strength parameter $\alpha Z$ ($\alpha$ is the fine-structure constant and $Z$ is the nuclear charge). In the present work, we employ the related formalism known as the Furry picture of QED Furry (1951). We note that the corresponding approach is relevant not only for highly charged ions Mohr et al. (1998); Yerokhin et al. (2003); Sapirstein and Cheng (2008); Glazov et al. (2011); Volotka et al. (2013); Shabaev et al. (2018); Indelicato (2019); Shabaev (2024), but also finds application in the case of light systems including the hydrogen atom, see, e.g., Refs. Jentschura et al. (1999); Yerokhin and Shabaev (2015b); Yerokhin et al. (2024); Mohr et al. (2025). An alternative, and in some sense complementary, nonrelativistic QED approach Caswell and Lepage (1986), primarily applicable to low-$Z$ systems, is beyond the scope of this paper. One of the leading QED corrections to energy levels originates from the one-electron self-energy (SE) diagram, shown in Fig. 1.
@@
However, high-precision evaluation of the corresponding contribution holds significance not only for studying the Lamb shift and transition energies in hydrogen-like ions and other few-electron systems Stöhlker et al. (2000); Gumberidze et al. (2005); Schweppe et al. (1991); Brandau et al. (2003); Beiersdorfer et al. (2005); Trassinelli et al. (2009, 2011); Loetzsch et al. (2024) (for related theory, see, e.g., Refs. Yerokhin et al. (2006); Kozhedub et al. (2008); Sapirstein and Cheng (2011); Yerokhin and Shabaev (2015a); Artemyev et al. (2005); Kozhedub et al. (2019)), but also for investigating a wide range of atomic properties. Accurate calculations of the SE contribution is a cornerstone in the ab initio treatment of the one-electron QED corrections to the $g$ factor Beier et al. (2000); Yerokhin et al. (2002), hyperfine structure Yerokhin and Shabaev (2001); Sapirstein and Cheng (2001); Yerokhin and Jentschura (2008, 2010a), quadratic Zeeman splitting Agababaev et al. (2025), nuclear-magnetic shielding Yerokhin et al. (2011, 2012), E1 Kozlov et al. (2024) and M1 Volotka et al. (2006) transition amplitudes. The SE contribution also plays an important role in the treatment of various many-electron QED effects, see, e.g., Yerokhin et al. (1997); Persson et al. (1996); Yerokhin et al. (1999); Volotka et al. (2009, 2014) and references therein. In all these applications it is often necessary to calculate off-diagonal matrix elements of the SE operator. However, the one-electron (diagonal) SE contribution is the simplest example that one can and should use to probe different methods and numerical approaches.
@@
Within the Furry-picture formalism, an issue of representing the bound-electron propagator in calculations naturally arises. Unlike the free-electron propagator, the propagator in the presence of external field generally does not have a closed-form expression. However, when the spherical symmetry is present, as is the case of atomic electrons, the partial-wave (PW) expansion helps to construct the propagator. With this in mind, we use two state-of-the-art approaches: the Green’s function (GF) method Mohr (1974b, a); Soff and Mohr (1988); Yerokhin and Maiorova (2020), and the finite-basis-set (FBS) method Johnson et al. (1988b); Tupitsyn and Loginov (2003); Kotochigova et al. (2007); Shabaev et al. (2004).
@@
Within the first approach, the bound-electron propagator is represented in terms of the Dirac-equation solutions bounded at infinity and at origin. For the point-nucleus Coulomb potential, the corresponding solutions are expressed analytically via the Whittaker functions Wichmann and Kroll (1956). For more realistic models of nuclear-charge distribution, e.g., the Fermi model, or other spherically symmetric potentials, e.g., some local screening potentials, these solutions can be found numerically. This approach enables one to perform calculations with a high numerical accuracy when large numerical cancellations occur or when the PW expansion does not converge rapidly. Note that the Green’s function, considered as a function of any of its radial arguments, is discontinuous, when the arguments coincide. Therefore, obtaining accurate results for the matrix elements of the Green’s function requires special care.
@@
Within the second approach, the electron Green’s function is represented using the FBS’s. The basis sets for the Dirac equation can be constructed, e.g., from the B splines Sapirstein and Johnson (1996); Johnson et al. (1988a); Shabaev et al. (2004), Sturm orbitals Tupitsyn and Loginov (2003); Kotochigova et al. (2007); Ivanov et al. (2024) or Gaussians Boys and Egerton (1950); Ferenc et al. (2025). The corresponding numerical procedures can easily incorporate any spherically symmetric potential. Moreover, for systems that do not possess spherical symmetry, the FBSs can also be readily prepared E. B. Rozenbaum, D. A. Glazov, V. M. Shabaev, K. E. Sosnova, and D. A. Telnov (2014); A. N. Artemyev and A. Surzhykov (2015); 1. The advantage of this method is that it allows one to easily separate out and exclude, if necessary, the contribution of a specific bound state to the Green’s function. Note also that this method provides an approximation to the Green’s function, which is a continuous function of the radial arguments. Nevertheless, the basis-set method has some important drawbacks when compared to the GF approach. Because of an additional parameter, the number of basis functions $N$, the obtained results should be extrapolated to $N\rightarrow\infty$. Typically, this sets a limit on the computational accuracy. Furthermore, the number of partial waves, usually accessible in practical calculations by means of the FBS method, is limited and smaller than that in the Green’s function method Yerokhin and Maiorova (2020). In this paper, we use the FBS method with the basis obtained from B splines within the dual-kinetic-balance (DKB) approach Shabaev et al. (2004). The latter method eliminates the so-called spurious states and establishes the correct asymptotics of wave functions in the non-relativistic limit. The effectiveness of this basis for calculating the SE correction has been demonstrated, e.g., in Ref. Shabaev et al. (2004) for the Coulomb potential and in Ref. Glazov et al. (2006) for local screening potentials.
@@
The SE diagram contains ultraviolet (UV) divergences. In the present work, these divergences are renormalized using the conventional approach based on the potential expansion of the bound-electron propagator Snyderman (1991). Within this expansion, three contributions naturally arise: zero-, one-, and many-potential terms. Only the zero- and one-potential terms, which correspond to the first and second terms of the potential expansion, are UV divergent. They are treated separately in momentum space together with the mass counterterm. The remaining many-potential term is considered in coordinate space using the PW expansion. The truncation of this expansion and subsequent estimation of the residual is generally the main source of numerical uncertainty of the results.
@@
The total SE contribution is gauge invariant, unlike the individual terms of its potential expansion. It is therefore intriguing to compare the magnitude of these terms in different gauges of the photon propagator. Starting from the pioneering works Mohr (1974b, a), the SE correction is calculated predominantly in the Feynman gauge, see, e.g., Refs. Mohr (1974a, 1982); Yerokhin and Shabaev (1999); Jentschura et al. (2001); Jentschura and Mohr (2004); Yerokhin and Harman (2017) and references therein. In the Coulomb gauge, similar calculations were carried out in Refs. Hedendahl and Holmberg (2012); Yerokhin et al. (2025a). These works suggest that the many-potential contribution is significantly smaller in the Coulomb gauge. However, a smaller absolute value does not necessarily imply better convergence of the PW expansion. Therefore, a thorough comparative analysis of the many-potential contributions in both gauges is in demand. Note that other covariant gauges, aside from the Feynman one, can also be applied to the SE calculations. For example, the Fried-Yennie gauge Adkins (1993) was also considered in Ref. Yerokhin et al. (2025a). In the present work, however, we restrict our consideration only to the Feynman and Coulomb gauges.
@@
Another important issue in the SE calculations is to somehow accelerate the convergence of PW expansions mentioned above. It is reasonable to assume that the poor convergence of the many-potential contribution is due to the leading term of its potential expansion. Therefore, the natural first step is to additionally subtract the two-potential term, which should improve the convergence of the remainder. The subtracted term has to be evaluated separately with high precision within some alternative method. In the following, we will refer to this approach as the two-potential scheme. It should be noted, however, that the direct calculation of the two-potential term in momentum space in a closed form, i.e., not resorting to the PW expansion, turns out to be a challenging problem Yerokhin and Jentschura (2008, 2010b).
@@
Another convergence-acceleration approach, hereafter refered to as the Sapirstein-Cheng (SC) scheme, was proposed recently in Ref. Sapirstein and Cheng (2023). Within the framework of this promising scheme, the quasi-two-potential contribution, which is an approximation to the two-potential term, is subtracted from the many-potential term instead. In contrast to the exact two-potential term, its quasi-two-potential counterpart can be easily calculated in momentum space. For the point-nucleus case, calculations of the SE correction using this acceleration scheme have been performed in Ref. Yerokhin et al. (2025a) for different gauges. However, the case of an extended nucleus is also worth examining.
@@
Yet another convergence-acceleration approach was developed earlier for the SE contribution to the Lamb shift in Ref. Yerokhin et al. (2005). We refer to it as the Yerokhin, Pachucki, and Shabaev (YPS) scheme. This scheme also improves the convergence by subtracting an approximation to the many-potential contribution, that can be calculated in a closed form. The YPS scheme was initially developed for the calculations in the Feynman gauge, and was later generalized to the Coulomb gauge in Ref. Yerokhin et al. (2025a). Judging by the results of the latter paper, it is an extremely powerful scheme, however, it is difficult to generalize it to more complex diagrams, unlike other convergence-acceleration approaches. For instance, the two-potential scheme has recently been applied in the QED calculations of the quadratic Zeeman effect Agababaev et al. (2025), and some modifications of the SC scheme have been applied to the calculations of the two-electron SE Malyshev et al. (2024) and two-loop SE Yerokhin et al. (2024, 2025b) contributions. For this reason, we do not consider the YPS scheme in the current study.
@@
Thus we thoroughly examine the application of several different schemes to the SE calculations in both the Coulomb and Feynman gauges. We conduct an independent of Ref. Yerokhin et al. (2025a) analysis, consider in details the issue of the PW-expansion convergence, evaluate the SE correction for finite-size nuclei, and collect the necessary formulas in a convenient form that allow one to make the corresponding calculations.
@@
Throughout this article the relativistic units ($\hbar=c=1$) and the Heaviside charge unit ($\alpha=e^{2}/(4\pi),e<0$) are used. We use roman style ($\rm{p}$) for four-vectors, bold style ($\bf{p}$) for space vectors, and italic style ($p$) for scalars.
@@ II.1 Basic formalism
Within the framework of the Furry picture, the description of a bound electron starts with the Dirac equation:
$\left[-i(\boldsymbol{\alpha}\cdot\boldsymbol{\nabla})+\beta m_{e}+V(\mathbf{r})\right]|a\rangle=\varepsilon_{a}|a\rangle,$ (1)
where $|a\rangle$ and $\varepsilon_{a}$ are the Dirac wave function and relativistic energy, $\boldsymbol{\alpha}$ and $\beta$ are the Dirac matrices, $V(\mathbf{r})$ is the binding potential, which is a function of the position vector $\mathbf{r}$. In the following, we consider only hydrogen-like ions where the binding potential is spherically symmetric, $V(\mathbf{r})=V(r)$, $r=|\mathbf{r}|$. Then, the solution of Eq. (1) can be represented in the form:
@@
Figure 1: Self-energy diagram with the related mass counterterm. The double line indicates the electron propagator in the external field of the nucleus, the wavy line denotes the photon propagator, and the cross stands for the mass counterterm.
The energy shift $\Delta\varepsilon_{a}$ of the bound state $|a\rangle$ due to the first-order self-energy correction, which is graphically represented in Fig. 1, is given by the real part of the following expression Mohr et al. (1998):
@@
where $\mathbf{r}_{12}=\mathbf{r}_{1}-\mathbf{r}_{2}$, $D_{\mu\nu}(\omega,\mathbf{r}_{12})$ is the photon propagator, $\alpha^{\mu}=(1,{\boldsymbol{\alpha}})$, $G(\varepsilon,\mathbf{r}_{1},\mathbf{r}_{2})$ is the bound-electron Green’s function, and $\delta m$ is the mass counterterm. Let us define the gauge-dependent operator $I(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ according to
@@
In the Feynman and Coulomb gauges, this operator takes the forms $I_{\text{F}}$ and $I_{\text{C}}$, respectively,
$I_{\text{F}}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\alpha\left[1-({\boldsymbol{\alpha}}_{1}\cdot{\boldsymbol{\alpha}}_{2})\right]\frac{e^{i\hat{\omega}r_{12}}}{r_{12}},$ (5)
$I_{\text{C}}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\alpha\left[\frac{1}{r_{12}}-({\boldsymbol{\alpha}}_{1}\cdot{\boldsymbol{\alpha}}_{2})\frac{e^{i\hat{\omega}r_{12}}}{r_{12}}+(\boldsymbol{\alpha}_{1}\cdot\boldsymbol{\nabla}_{1})(\boldsymbol{\alpha}_{2}\cdot\boldsymbol{\nabla}_{2})\frac{e^{i\hat{\omega}r_{12}}-1}{\omega^{2}r_{12}}\right],$ (6)
where $r_{12}=|\mathbf{r}_{12}|$, $\hat{\omega}=\sqrt{\omega^{2}+i0}$, and the branch of the square root is fixed with the condition $\text{Im \,}\sqrt{\omega^{2}+i0}>0$.
The bound-electron propagator $G$ can be expressed using the spectral representation,
$G(\omega,\mathbf{r}_{1},\mathbf{r}_{2})=\sum_{n}\frac{\psi_{n}(\mathbf{r}_{1})\psi_{n}^{\dagger}(\mathbf{r}_{2})}{\omega-\varepsilon_{n}(1-i0)},$ (7)
where the summation runs over the complete Dirac spectrum. In shortened form, the formula (7) can be written as follows:
$G(\omega)=\sum_{n}\frac{|n\rangle\langle n|}{\omega-\varepsilon_{n}^{-}},$ (8)
where the notation $\varepsilon_{n}^{-}=\varepsilon_{n}(1-i0)$ was introduced. Then, the energy shift (3) can be expressed as:
@@
where $\gamma^{0}\equiv\beta$ and the diagonal matrix element of the one-loop self-energy operator $\Sigma(E)$ is given by:
$\langle a|\gamma^{0}\Sigma(E)|a\rangle=\frac{i}{2\pi}\int_{-\infty}^{\infty}d\omega\sum_{n}\frac{\langle an|I(\omega)|na\rangle}{E-\omega-\varepsilon_{n}^{-}}.$ (10)
The expression (10) contains the UV divergences. To properly treat them, we expand the bound-electron propagator $G$ in powers of the binding potential $V$ Blundell and Snyderman (1991). The potential expansion can be conveniently written as:
@@
where the states $|n\rangle$ correspond to the bound-electron spectrum, the states $|f\rangle$, $|f_{1}\rangle$, and $|f_{2}\rangle$ correspond to the free-electron spectrum, and the last term is expressed via the bound-electron Green’s function $G(\omega)$ itself and encompasses an infinite number of terms corresponding to the two or more interactions with the nucleus. As before, all summations in Eq. (11), are carried out over the complete spectra for the corresponding Dirac Hamiltonians.
Figure 2: Potential expansion of the self-energy diagram. The dashed line ended by a rhombus denotes the interaction with the binding potential. The single line denotes the free-electron propagator. The mass counterterm is not shown.
Substituting the potential expansion (11) into Eq. (10) results in the SE diagram being split into three contributions, as shown graphically in Fig. 2. These contributions are commonly referred to as the zero-, one-, and many-potential terms. The zero- and one-potential terms diverge, whereas the many-potential term is UV finite. The UV-divergent terms have to be renormalized along with the mass counterterm. For this aim, they are considered in momentum space, and the dimensional regularization is applied to isolate and eliminate the UV divergences. This renormalization procedure yields the UV-finite contributions $\Delta\varepsilon_{a}^{\text{0p}}$ and $\Delta\varepsilon_{a}^{\text{1p}}$ to the Lamb shift. The many-potential term $\Delta\varepsilon_{a}^{\text{Mp}}$ is calculated in the coordinate space. Finally, the SE correction to energy levels is expressed as the gauge-invariant sum of three contributions:
@@
In the following subsections we consider in detail these terms and some specific methods to improve the convergence of the partial-wave expansion in the many-potential part.